OpenStudy (johnnydicamillo):

Given that f(x)=x2+7 and g(x)=x–1, calculate fog

OpenStudy (johnnydicamillo):

and find the domain

OpenStudy (johnnydicamillo):

I got \[x^{2}-2x+8\]

OpenStudy (anonymous):

i got {x^2+7 = f(x), x = g(x)+1}

OpenStudy (imstuck):

Work from the inside out. Take the g function and put the whole thing into the f function where you see an x. The g function is x-1. So put x-1 into the x place in the f function. Like this: (x-1)^2 + 7. You won't find a numerical answer because you don't have a value for x, but you can find the composite function.